Del Monty will receive the following payments at the end of the next three years: $24,000, $27,000, and $29,000. Then from the end of the 4th year through the end of the 10th year, he will receive an annuity of $30,000 per year.
 

At a discount rate of 16 percent, what is the present value of all three future benefits? Use Appendix B and Appendix D for an approximate answer, but calculate your final answer using the formula and financial calculator methods. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
Answer #1

We know, Present value of inflows = cash inflow * Present value of discounting factor (rate%, time period)


so,

= 24000/1.16+27000/1.16^2+29000/1.16^3+30000/1.16^4+30000/1.16^5+........+30000/1.16^10

= (24000*0.862068965)+(27000*0.743162901)+(29000*0.640657673)+(30000[1/1.16^4+1/1.16^5+1/1.16^6+....+1/1.16^10]

= (24000*0.862068965)+(27000*0.743162901)+(29000*0.640657673)+(30000*2.587337938)

=$136,954.26(Approx)

ps: intermediate calculations have not been rounded off.


answered by: gavin